6. Ship A is travelling south at the rate of 2 km/hr, at the instant that ship B, which is 32


miles south of ship A, is travelling east at rate of 4 km/hr.



a) Are they separating or approaching at the end of 2 hrs, and at what rate?



b) At what time are they nearest together?



c) What is their minimum distance apart?

1
Expert's answer
2021-12-22T15:53:00-0500

a)

initial distance between A and B is 32 km

distance between A and B after 2 hours:

"\sqrt{(32-2\cdot2)^2+(4\cdot2)^2}=29.12" km < 32 km

so, A and B are approaching at the end of 2 hrs


rate of approaching at the end of 2 hrs:


"d'(2)=\frac{16\cdot2-2(32-2\cdot2)}{\sqrt{(32-2\cdot2)^2+(4\cdot2)^2}}=-0.82" km/h


b)

distance between A and B as function of time:

"d(t)=\sqrt{(32-2t)^2+(4t)^2}"


"d'(t)=\frac{16t-2(32-2t)}{\sqrt{(32-2t)^2+(4t)^2}}=0"


"16t-2(32-2t)=0"

"t=64\/20=3.2" hours


c)

minimum distance apart:

"d(3.2)=\sqrt{(32-2\cdot3.2)^2+(4\cdot3.2)^2}=28.62" km